Home CMS Production Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS June 2009  Not logged in ELOG logo
Message ID: 63     Entry time: Sat Aug 22 10:28:50 2009
Author: Thomas Bergauer 
Project: BELLE 
Subject: APVDAQ + telescope 40 MHz hit mode (cooling?) 
Run Number: run017 
Events: 75k 
StartTime: 22.08.2009 10:26:28 
EndTime: 22.08.2009 11:16:12 

 HV=80V, 40 MHz, beam on, 120 GeV

6 samples, hit mode, hitcut=4
chiller set to 13 deg.
 

f=40 MHz

config file: cern09_belle_hit_multi6.cfg

 

there was probably no water flow throught the pipe of the origami module during this run!

ELOG V3.1.4-395e101